an airplane is heading 35 degrees counter-clockwise and it turns towards 102 degrees counter-clockwise with a turning-radius of 2700m.
how long will the turn last if the speed of the airplane is 360km/h ?

Answer by stanbon(75887)

an airplane is heading 35 degrees counter-clockwise and it turns towards 102 degrees counter-clockwise with a turning-radius of 2700m = 27km.
how long will the turn last if the speed of the airplane is 360km/h ?
---
The angle turn is 102-35 = 67 degrees.
---
67 degrees = 67(pi/180 = 0.3722(pi) radians
---
Arc length = r*theta = 27(0.3722(pi)) km
----
Time = [27*0.3722(pi)]/[360km/h] = 0.0877 hrs = 5.26 minutes
===============================================================
Cheers,
Stan H.
============================